REINVESTIGATION OF THE GROUND AND FIRST TORSIONAL STATES OF METHYLFORMATE

M.. CARVAJAL, Departamento de Fisica Aplicada, Facultad de Ciencias Experimentales, Universitad de Huelva, 21071 Huelva, Spain; F.. WILLAERT, J.. DEMAISON, Laboratoire de Physique des Lasers, Atomes et Molécules, CNRS et Université de Lille I, F-59655, Villeneuve d’Ascq Cédex, France; \underline{I. KLEINER}, Laboratoire Interuniversitaire des Systèmes Atmosphériques, CNRS et Universités Paris 7 et Paris 12, 61 av. Général de Gaulle, 94010, Créteil, France.

We have reinvestigated the laboratory spectrum for the methylformate molecule involving both the ground and first torsional states. We have fitted within almost experimental accuracy a data set for HCOOCH_{3} consisting of 3496 v_{t}=0 and 774 v_{t}=1 microwave lines, using the so-called ”rho axis method” (RAM) and a model extended to include perturbation terms through eight order. The previously published v_{t}=0 and v_{t}=1 microwave lines covering J values up to 43 in the ground and up to 18 in the first excited states have been extended by new measurements from Lille with maximum values of J = 62. The final fit requires only 49 parameters to achieve a weigthed standard deviation for the whole fit of 1.43 for a total of 4270 lines. This result represents an improvement over the previous fit which achieved a standard deviation of 1.96 for 3862 lines using 69 parameters. A calculation of the linestrengths of torsion-rotation transition up to J=60 needed for the astronomical survey is also presented.